# A Nonparametric Bayesian Solution of the Empirical Stochastic Inverse Problem

### v1

@@ Appendix A Proofs
For all $\lambda\in Q^{-1}(q_{0})$, there exists a $N(\lambda,r)$ such that there exists $\{i_{1}$, $\dots$, $i_{m}\}$ $\subset$ $\{1,\dots,n\}$ with $J_{Q}^{(m)}=\left(J_{Q}^{(i_{1})}\,J_{Q}^{(i_{2})}\,\cdots J_{Q}^{(i_{m})}\right)$ being full rank on $Q^{-1}(q_{0})$ in $N(\lambda,r)$. The boundary of $N(\lambda,r)\cap\Lambda$, which is the union of part of $\partial\Lambda$ and part of $\partial N(\lambda,r)$ is piecewise smooth. Suppose $\partial\left(N(\lambda,r)\cap\Lambda\right)$ is locally determined by $B^{\prime}(\lambda)=0$. When $r$ is sufficiently small, $\big(J_{Q}\;J_{B^{\prime}}\big)^{\top}$ is full rank on $\partial\left(N(\lambda,r)\cap\Lambda\right)\cap Q^{-1}(q_{0})$ by continuity of $J_{Q}$. The collection $\{N(\lambda,r)\}$ admits a finite sub cover $\{N(\lambda_{k},r_{k})\}_{k=1}^{K}$ of $\Lambda\cap Q^{-1}(q_{0})$. Taking $r$ sufficiently small, the collection covers $\Lambda\cap Q^{-1}(q)$, for $q$ in $N(q_{0},r)$. Define $N_{k}=N(\lambda_{k},r_{k})\setminus\left(\bigcup_{j=1}^{k-1}N(\lambda_{j},r_{j})\right)$, the boundary of $N_{k}\cap\Lambda$ is piecewise smooth. Suppose $\partial\left(N_{k}\cap\Lambda\right)$ is locally determined by $B_{k}(\lambda)=0$, then, $\big(J_{Q}\;J_{B_{k}}\big)^{\top}$ is full rank on $\partial\left(N_{k}\cap\Lambda\right)$, and
$\tilde{\rho}_{\mathcal{D}}(q)$ $=\int_{Q^{-1}(q)\cap\Lambda}\frac{1}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds=\sum_{k=1}^{K}\int_{Q^{-1}(q)\cap\left(N_{k}\cap\Lambda\right)}\frac{1}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds.$
@@
$|\tilde{\rho}_{\mathcal{D}}(q)-\tilde{\rho}(q_{0})|\leq\sum_{k=1}^{K}\left|\int_{Q^{-1}(q)\cap\left(N_{k}\cap\Lambda\right)}\frac{1}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds-\int_{Q^{-1}(q_{0})\cap\left(N_{k}\cap\Lambda\right)}\frac{1}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds\right|$
Theorem A.3 implies that $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous at $q_{0}$. Since $\int_{Q^{-1}(q)\cap\left(N_{k}\cap\Lambda\right)}\frac{1}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds$ is continuous in a neighborhood of $q_{0}$, $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous in a neighborhood of $q_{0}$. ∎
@@
We begin by developing and analyzing numerical solution method for the SIP built on simple function approximations of the integrals in the disintegration (2.5) associated with sequences of partitions of $\Lambda$ and $\mathcal{D}$ where random sampling is used to create the partitions. This forms the theoretical basis for the practical sampling-based method used in practice. The approach is heavily influenced by the theory of stochastic geometry and the approximation of sets [67, 45, 56].
@@
The foundation of the approximation of the solution of the SIP are sequences of simple functions corresponding to sequences of partitions of $\mathcal{D}$ and $\Lambda$. For a sequence of integers $0<M_{1}<M_{2}<\cdots$, we consider partitions $\{\mathcal{I}_{M_{\ell}}\}_{\ell=1}^{\infty}$ of $\mathcal{D}$, with $\mathcal{I}_{M_{\ell}}=\big\{I_{i}^{(\ell)}\big\}_{i=1}^{M_{\ell}}$ and associated points $\big\{\{d^{(\ell)}_{i}\}_{i=1}^{M_{\ell}}\big\}_{\ell=1}^{\infty}$, with $d^{(\ell)}_{i}\in I_{i}^{(\ell)}$ for all $i,\ell$. Likewise, for a sequence of integers $0<N_{1}<N_{2}<\cdots$, we consider partitions $\{\mathcal{T}_{N_{\jmath}}\}_{\jmath=1}^{\infty}$ of $\Lambda$, with $\mathcal{T}_{N_{\jmath}}=\big\{B_{i}^{(\jmath)}\big\}_{i=1}^{N_{\jmath}}$ and associated points $\big\{\{\lambda^{(\jmath)}_{i}\}_{i=1}^{N_{\jmath}}\big\}_{\jmath=1}^{\infty}$, with $\lambda^{(\jmath)}_{i}\in B_{i}^{(\jmath)}$ for all $i,\jmath$.
@@
The formal approximation treated in Theorem A.6 is impractical in general because computing the measures of cells in the partitions $\{\mathcal{I}_{M_{\ell}}\}_{\ell=1}^{\infty}$ and $\{\mathcal{T}_{N_{\jmath}}\}_{\jmath=1}^{\infty}$ is computationally untenable. However, the result clarifies the need for the assumptions on the samples used in the numerical solution.
@@
as an approximation of ${\rho}_{\Lambda}$. By definition, ${\rho}_{\Lambda,M_{\ell}}(\lambda)=\rho^{\prime}_{\mathcal{D},M_{\ell}}(Q(\lambda))$. By Theorem A.5, $\rho^{\prime}_{\mathcal{D},M_{\ell}}$ converges to $\rho_{\mathcal{D}}^{\prime}=\frac{\rho_{\mathcal{D}}}{\tilde{\rho}_{\mathcal{D}}}$, $\tilde{\mu}_{\mathcal{D}}$-a.e., so ${\rho}_{\Lambda,M_{\ell}}\to{\rho}_{\Lambda}$, $\mu_{\Lambda}$-a.e., where ${\rho}_{\Lambda}(\lambda)=\frac{\rho_{\mathcal{D}}(Q(\lambda))}{\tilde{\rho}_{\mathcal{D}}(Q(\lambda))}$ is the density for the uniform prior solution.
@@
$\mathbb{E}\left(\frac{\sum_{k=1}^{N_{\jmath}}\chi_{\lambda_{k}}(Q^{-1}(I_{i}^{(K)})\cap A)}{\sum_{j=1}^{N_{\jmath}}\chi_{\lambda_{j}}(Q^{-1}(I_{i}^{(K)}))}\right)$ $=\sum_{k=1}^{N_{\jmath}}E\,\left(\frac{\chi_{\lambda_{k}}(Q^{-1}(I_{i}^{(K)})\cap A)}{\sum_{j=1}^{N_{\jmath}}\chi_{\lambda_{j}}(Q^{-1}(I_{i}^{(K)}))}\right)$

### v3

@@ Appendix A Proofs
For all $\lambda\in Q^{-1}(q_{0})$, there exists a $N(\lambda,r)$ such that there exists $\{i_{1}$, $\dots$, $i_{m}\}$ $\subset$ $\{1,\dots,n\}$ with $J_{Q}^{(m)}=\left(J_{Q}^{(i_{1})}\,J_{Q}^{(i_{2})}\,\cdots J_{Q}^{(i_{m})}\right)$ being full rank on $Q^{-1}(q_{0})$ in $N(\lambda,r)$. The boundary of $N(\lambda,r)\cap\Lambda$, which is the union of part of $\partial\Lambda$ and part of $\partial N(\lambda,r)$ is piecewise smooth. Therefore, there is a continuously differentiable map $B^{\prime}$ such that $\partial\left(N(\lambda,r)\cap\Lambda\right)$ is locally determined by $B^{\prime}(\lambda)=0$. When $r$ is sufficiently small, $\big(J_{Q}$ $J_{B^{\prime}}\big)^{\top}$ is full rank on $\partial\left(N(\lambda,r)\cap\Lambda\right)\cap Q^{-1}(q_{0})$ by continuity of $J_{Q}$. The collection $\{N(\lambda,r)\}$ admits a finite sub cover $\{N(\lambda_{k},r_{k})\}_{k=1}^{K}$ of $\Lambda\cap Q^{-1}(q_{0})$. Taking $r$ sufficiently small, the collection covers $\Lambda\cap Q^{-1}(q)$, for $q$ in $N(q_{0},r)$. Define $N_{k}=N(\lambda_{k},r_{k})\setminus\left(\bigcup_{j=1}^{k-1}N(\lambda_{j},r_{j})\right)$, the boundary of $N_{k}\cap\Lambda$ is piecewise smooth. We assume there exists $B_{k}(\lambda)$ such that $\partial\left(N_{k}\cap\Lambda\right)$ is determined locally by $B_{k}(\lambda)=0$, $k=1,\cdots,K$. Then, $\big(J_{Q}\;J_{B_{k}}\big)^{\top}$ is full rank on $\partial\left(N_{k}\cap\Lambda\right)$ and
$\tilde{\rho}_{p,\mathcal{D}}(q)$ $=\int_{Q^{-1}(q)\cap\Lambda}\frac{\rho_{p}}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds=\sum_{k=1}^{K}\int_{Q^{-1}(q)\cap\left(N_{k}\cap\Lambda\right)}\frac{\rho_{p}}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds.$
@@
$|\tilde{\rho}_{p,\mathcal{D}}(q)-\tilde{\rho}_{p,\mathcal{D}}(q_{0})|\\ \leq\sum_{k=1}^{K}\left|\int_{Q^{-1}(q)\cap\left(N_{k}\cap\Lambda\right)}\frac{\rho_{p}}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds-\int_{Q^{-1}(q_{0})\cap\left(N_{k}\cap\Lambda\right)}\frac{\rho_{p}}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds\right|$
Theorem A.3 implies that $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous at $q_{0}$. Since $\int_{Q^{-1}(q)\cap\left(N_{k}\cap\Lambda\right)}\frac{\rho_{p}}{\sqrt{\det(J_{Q}J_{Q}^{\top})}}\,ds$ is continuous in a neighborhood of $q_{0}$, $\tilde{\rho}_{p,\mathcal{D}}(q)$ is continuous in a neighborhood of $q_{0}$. ∎
@@
We begin by developing and analyzing an estimator built on simple function approximations of the integrals in the disintegration (2.5) associated with sequences of partitions of $\Lambda$ and $\mathcal{D}$ where random sampling is used to create the partitions. This forms the theoretical basis for the practical sampling-based method used in practice. The approach is heavily influenced by the theory of stochastic geometry and the approximation of sets [68, 46, 57].
@@
The foundation of the approximation of the posterior are sequences of simple functions corresponding to sequences of partitions of $\mathcal{D}$ and $\Lambda$. For a sequence of integers $0<M_{1}<M_{2}<\cdots$, we consider partitions $\{\mathcal{I}_{M_{\ell}}\}_{\ell=1}^{\infty}$ of $\mathcal{D}$, with $\mathcal{I}_{M_{\ell}}=\big\{I_{i}^{(\ell)}\big\}_{i=1}^{M_{\ell}}$ and associated points $\big\{\{d^{(\ell)}_{i}\}_{i=1}^{M_{\ell}}\big\}_{\ell=1}^{\infty}$, with $d^{(\ell)}_{i}\in I_{i}^{(\ell)}$ for all $i,\ell$. Likewise, for a sequence of integers $0<N_{1}<N_{2}<\cdots$, we consider partitions $\{\mathcal{T}_{N_{\jmath}}\}_{\jmath=1}^{\infty}$ of $\Lambda$, with $\mathcal{T}_{N_{\jmath}}=\big\{B_{i}^{(\jmath)}\big\}_{i=1}^{N_{\jmath}}$ and associated points $\big\{\{\lambda^{(\jmath)}_{i}\}_{i=1}^{N_{\jmath}}\big\}_{\jmath=1}^{\infty}$, with $\lambda^{(\jmath)}_{i}\in B_{i}^{(\jmath)}$ for all $i,\jmath$.
@@
The formal approximation treated in Theorem A.6 is impractical in general because computing the measures of cells in the partitions $\{\mathcal{I}_{M_{\ell}}\}_{\ell=1}^{\infty}$ and $\{\mathcal{T}_{N_{\jmath}}\}_{\jmath=1}^{\infty}$ is computationally untenable. However, the result clarifies the need for the assumptions on the samples used for the estimation.
@@
as an approximation of ${\rho}_{\Lambda}$. By definition, ${\rho}_{\Lambda,M_{\ell}}(\lambda)=\rho^{\prime}_{\mathcal{D},M_{\ell}}(Q(\lambda))$. By Theorem A.5, $\rho^{\prime}_{\mathcal{D},M_{\ell}}$ converges to $\rho_{\mathcal{D}}^{\prime}=\frac{\rho_{\mathcal{D}}}{\tilde{\rho}_{\mathcal{D}}}$, $\tilde{\mu}_{\mathcal{D}}$-a.e., so ${\rho}_{\Lambda,M_{\ell}}\to{\rho}_{\Lambda}$, $\mu_{\Lambda}$-a.e., where ${\rho}_{\Lambda}(\lambda)=\frac{\rho_{\mathcal{D}}(Q(\lambda))}{\tilde{\rho}_{\mathcal{D}}(Q(\lambda))}$ is the density for the uniform prior posterior.
@@
$E\left(\frac{\sum_{k=1}^{N_{\jmath}}\chi_{\lambda_{k}}(Q^{-1}(I_{i}^{(K)})\cap A)}{\sum_{j=1}^{N_{\jmath}}\chi_{\lambda_{j}}(Q^{-1}(I_{i}^{(K)}))}\right)$ $=\sum_{k=1}^{N_{\jmath}}E\,\left(\frac{\chi_{\lambda_{k}}(Q^{-1}(I_{i}^{(K)})\cap A)}{\sum_{j=1}^{N_{\jmath}}\chi_{\lambda_{j}}(Q^{-1}(I_{i}^{(K)}))}\right)$